# Preserving Membership Privacy in Low-Rank Adaptation for Latent Diffusion Models

### v2

@@ 4.2 Ablation Study
Table 7 in Appendix 6.6 displays the performance of SMP-LoRA with different learning rates $\eta_{2}\in\{1e-4,1e-5,1e-6\}$ on the Pokemon dataset. SMP-LoRA consistently preserves membership privacy across all tested learning rates.
@@
Table 8 in Appendix 6.6 shows the performance of SMP-LoRA with different rank $r\in\{128,64,32,16,8\}$ on the Pokemon dataset. The performance of SMP-LoRA is not significantly affected by the LoRA’s rank $r$.
@@
In Table 9 (Appendix 6.6), we present the performance of SMP-LoRA and its extension to the full fine-tuning and DreamBooth (Ruiz et al., 2023) methods on the Pokemon dataset. Our membership-privacy-preserving method continues to effectively protect membership privacy when applied to these methods, highlighting its potential applicability across different adaptation methods.
@@
In Table 10 (Appendix 6.6), we report the performance of SMP-LoRA and MP-LoRA enhanced with gradient clipping and normalization techniques on the Pokemon dataset. These traditional techniques for stabilizing the gradient cannot address the unstable optimization issue in MP-LoRA.
@@
In Table 11 (Appendix 6.6), we display the attack performance of the black-box MI attack (Wu et al., 2022) and the white-box gradient-based MI attack (Pang et al., 2023) on LoRA and SMP-LoRA. Compared to LoRA, SMP-LoRA, specifically designed to defend against white-box loss-based MI attacks, consistently provides enhanced membership privacy protection when facing MI attacks in different settings. Implementation details for the black-box and white-box gradient-based MI attacks are available in Appendix 6.7.
5 Conclusion

### v3

@@ Appendix F Detailed Ablation Study
Table 7 displays the performance of SMP-LoRA with different learning rate $\eta_{2}\in\{1e-4,1e-5,1e-6\}$ on the Pokemon dataset. We observe an increase in the FID scores as the learning rate $\eta_{2}$ decreases. This phenomenon might be due to the lower learning rate, which results in the model underfitting after 400 training epochs. Notably, across all tested learning rates, SMP-LoRA consistently preserves membership privacy, effectively defending against MI attacks.
@@
Table 8 shows the performance of SMP-LoRA with different rank $r\in\{128,64,32,16,8\}$ on the Pokemon dataset. We observe that the LoRA’s rank $r$ does not significantly affect the performance of SMP-LoRA.
@@
Table 9 presents the performance of SMP-LoRA and its extension to the full fine-tuning and DreamBooth (Ruiz et al. 2023) methods on the Pokemon dataset. Both the full fine-tuning and DreamBooth methods are sensitive to the learning rate. Accordingly, the learning rate $\eta_{2}$ for both methods was set at 5e-6, while other hyperparameters remained consistent with our experimental setup in Appendix E. In Table 9, we can observe that our membership-privacy-preserving method effectively protects membership privacy when applied to both methods. This underscores the potential applicability of our membership-privacy-preserving method across different adaptation methods.
Table 8: The effect of LoRA’s rank $r$ for SMP-LoRA on the Pokemon dataset.
$r$ FID $\downarrow$ ASR (%) $\downarrow$ $\frac{\left|\mathrm{AUC}-0.5\right|}{0.5}$downarrow
128 0.337 56.1 0.02
64 0.274 50.3 0.16
32 0.512 54.0 0.14
16 0.541 53.1 0.16
8 0.661 48.9 0.08
Table 9: Performance of SMP-LoRA and its extension to the full fine-tuning and DreamBooth (Ruiz et al. 2023) methods on the Pokemon dataset.
Method FID $\downarrow$ ASR (%) $\downarrow$ $\frac{\left|\mathrm{AUC}-0.5\right|}{0.5}$downarrow
LoRA 0.20$\pm$0.04 82.27$\pm$4.38 0.73$\pm$0.09
SMP-LoRA 0.32$\pm$0.07 51.97$\pm$1.20 0.14$\pm$0.02
Full Fine-tuning 0.176 80.4 0.66
SMP Full Fine-tuning 1.05 54.5 0.26
DreamBooth 0.260 80.6 0.70
SMP-DreamBooth 0.748 56.4 0.10
Table 10: Performance of SMP-LoRA and MP-LoRA enhanced with gradient clipping and normalization techniques on the Pokemon dataset.
Method FID $\downarrow$ ASR (%) $\downarrow$ $\frac{\left|\mathrm{AUC}-0.5\right|}{0.5}$downarrow
MP-LoRA + Gradient Clipping 3.513 46.2 0.14
MP-LoRA + Gradient Normalization 2.390 52.4 0.06
SMP-LoRA 0.274 50.3 0.16
Table 11: Performance of LoRA and SMP-LoRA under the black-box (Wu et al. 2022) and the white-box gradient-based (Pang et al. 2023) MI attacks on the Pokemon dataset.
Attack Method ASR (%) $\downarrow$ $\frac{\left|\mathrm{AUC}-0.5\right|}{0.5}$downarrow
Black-box LoRA 72.4 0.34
SMP-LoRA 55.4 0.08
White-box Gradient-based LoRA 92.4 0.84
SMP-LoRA 63.5 0.28
Figure 4: Generated results on the CelebA__\__Small and CelebA__\__Large datasets. Each column of three images is generated using the same text prompt.
@@
In Table 10, we report the performance of SMP-LoRA and MP-LoRA enhanced with gradient clipping and normalization techniques on the Pokemon dataset. Notably, gradient clipping is a standard technique in training Latent Diffusion Models and has been employed in our experiments. In Table 10, the higher FID scores for gradient clipping (3.513) and gradient normalization (2.390) reflect poor image quality, indicating the failure in adaptation. These results demonstrate that such traditional techniques for stabilizing the gradient, as gradient clipping and normalization, cannot effectively address the unstable optimization issue in MP-LoRA.
@@
Table 11 displays the attack performance on LoRA and SMP-LoRA using the black-box MI attack (Wu et al. 2022) and the white-box gradient-based MI attack (Pang et al. 2023), which is the currently the most potent MI attack targeting DMs. Compared to LoRA, SMP-LoRA, specifically designed to defend against white-box loss-based MI attacks, exhibits lower ASR and better AUC, indicating that it can still preserve membership privacy to a certain extent when facing MI attacks in different settings. Until now, we have evaluated SMP-LoRA against the white-box loss-based MI attack and the currently strongest MI attack, the white-box gradient-based MI attack, rendering further comparisons with weaker attacks unnecessary, such as gray-box MI attacks (Duan et al. 2023; Kong et al. 2024; Fu et al. 2023). The black-box MI attack was implemented from the semantic-based Attack II-S proposed by Wu et al. (2022), and the white-box gradient-based MI attack was replicated from the GSA1 proposed by Pang et al. (2023). Implementation details for both MI attacks are available in Appendix G.
Appendix G Implementation Details
In this section, we detail the implementation of the black-box and white-box gradient-based MI attacks used in Section 4. The attack performance of both attacks on LoRA-adapted and SMP-LoRA-adapted LDMs is presented in Table 11.
For the black-box MI attack, we utilize the semantic-based Attack II-S proposed by Wu et al. (2022). This attack leverages the pre-trained BLIP model to extract embeddings for both an image and its corresponding text-generated image, and then conduct MI attacks based on the L2 distance between these two embeddings. Based on Wu et al. (2022)’s experiment setup, we instantiate the attack model as a 3-layer MLP with cross-entropy loss, optimized with Adam at a learning rate of 1e-4 over 200 training epochs.
For the white-box gradient-based MI attack, we replicate the GSA1 proposed by Pang et al. (2023). This attack involves uniformly sampling ten steps from the total diffusion steps, calculating the loss at each step, averaging these losses, and then performing backpropagation to obtain gradients. These gradients are then used to train an XGBoost model to infer membership and non-membership.
Appendix H Visualization of Generated Results
We provide more generated results of LoRA-adapted, MP-LoRA-adapted, and SMP-LoRA-adapted LDMs on the CelebA__\__Small and CelebA__\__Large datasets in Figure 4.
Appendix I Discussion
MP-LoRA vs. SMP-LoRA
In Table 1, it is evident that MP-LoRA significantly enhances membership privacy but compromises image generation capabilities, while SMP-LoRA effectively preserves membership privacy without impairing image quality. Moreover, in terms of ASR and TPR@$5\%$FPR metrics, SMP-LoRA sometimes outperforms MP-LoRA in preserving membership privacy. Based on these findings, we cannot assert that MP-LoRA provides better membership privacy protection compared to SMP-LoRA, as MP-LoRA is not fully optimized, evidenced by its complete loss of image generation capability. Therefore, SMP-LoRA is not just a balance within the utility-privacy Pareto Optimality.
Appendix J Limitations
Our SMP-LoRA method is primarily designed to defend against white-box loss-based MI attacks, yet there are other aspects of privacy vulnerability, such as Model Inversion Attack (Carlini et al. 2023). Besides, our method is currently applied solely to LoRA (tested on full fine-tuning and DreamBooth), and its extension to other adaptation methods, such as Textual Inversion (Gal et al. 2023) and Hypernetwork (NovelAI 2022), remains unexplored. Additionally, our selection of the coefficient value $\lambda$ relies on empirical validations, lacking a principled way.